OpenStudy (anonymous):

110kg of masses are on top of a 10 kg sled. it is pulled at a 37 degree angle, with respect to the ground. once in motion it will move at a horizontal, constant velocity. The coefficient of friction is .45. Determine the amount of force being applied along the angle?

OpenStudy (anonymous):

What part of this question are you stuck on? Have you drawn a free-body diagram? Do you know the equation for the force exerted by friction? Where, specifically, do you need help?

OpenStudy (anonymous):

Yes, I've drawn one, I'm not sure if it's taught this way in other places, but I divided the angled force into the x and y directions( T subscript x and T subscript y) and made a summation statement for both (Ty+Fn-Fg=0 and Tx-Ff=0) I'm assuming that the sum of Tx and Ty would give me the T I'm looking for, but I've only gotten as far as trying to substitute the two with numbers I was given. So far I've gotten Ty-1176=Fn and Tx-(.45*[Ty-1176]) But I don't know where to go from there.

OpenStudy (anonymous):

Oh for that second formula I meant Tx-(.45*[Ty-1176])=0, sorry!

OpenStudy (anonymous):

I think you are on the right track. I haven't checked your numbers, but your setup looks correct so far. Sum of the forces in each direction equals zero because there is no acceleration. So, you have two equations and three variables, which obviously won't work. You need a 3rd equation to be able to solve the system. Since you are given the angle at which the rope is being pulled, you know that T = Tx(cos(37)) + Ty(sin(37)). This is your 3rd equation that should allow you to solve the system for T. Does that make sense? I'll work through the rest of it to check the numbers.

OpenStudy (anonymous):

Ah, ok I see where you're going with this.. I'll go ahead and try to work with it, thank you very much!

OpenStudy (anonymous):

You are welcome. If this works for you, please mark it as answered. I'll finish working through the problem in case you have further questions.

OpenStudy (anonymous):

I mis-spoke above. What I meant to say for T is that T = Tx + Ty where Tx=T*cos(37) and Ty=T*sin(37). I think you understood what I meant, but I wrote it wrong. I worked the rest of the way through the problem and got 495 N. \[T=\mu*Fg/(\cos(37)+\mu*\sin(37)) \]